# Homogenized Equations for Isentropic Gas in a Pipe with Periodically-Varying Cross-Section

### v1

@@ 4.1 Numerical Discretization of Homogenized System
We discretize the full nonlinear homogenized equations, with the 3rd-order derivative terms written as in the stable formulation (3.9), found in 3:
$\underline{\rho}_{t}$ $=\alpha_{1}\underline{q}_{x}+\alpha_{2}\underline{q}_{xx}+\alpha_{3}\underline{q}\,\underline{\rho}_{x}+\alpha_{4}\underline{q}^{2}\underline{q}_{x}+\alpha_{5b}\underline{\rho}_{xxt}+\alpha_{6}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\alpha_{7}\underline{q}_{x}\underline{\rho}_{x}+\alpha_{8}\underline{q}\,\underline{\rho}_{xx},$ (4.1a)
$\underline{q}_{t}$ $=\beta_{1}\underline{\rho}_{x}+\beta_{2}\underline{q}\,\underline{q}_{x}+\beta_{3}\underline{\rho}\,\underline{\rho}_{x}+\beta_{4}\underline{\rho}_{xx}+\beta_{5}\underline{q}\,\underline{\rho}\,\underline{q}_{x}+\beta_{6}\underline{q}_{x}^{2}+\beta_{7}\underline{q}\,\underline{q}_{xx}+\beta_{8}\underline{q}^{2}\underline{\rho}_{x}$ (4.1b)
$\ \ +\beta_{9}\underline{\rho}_{x}^{2}+\beta_{10}\underline{\rho}\,\underline{\rho}_{xx}+\beta_{11b}\underline{q}_{xxt}.$
We first rewrite (4.1) in the form
@@
Here, $\mathcal{F}_{1}(\overline{\rho},\overline{q})$ and $\mathcal{F}_{2}(\overline{\rho},\overline{q})$ are the right-hand sides of the first and second equations in (4.1), respectively, excluding the $xxt$-derivative terms. We solve this system using a Fourier pseudospectral collocation method in space and a third-order strong stability preserving Runge-Kutta method [21] in time.
@@ 4.3.1 Piecewise-Constant Cross-section
In this section, we consider flow in a pipe with piecewise-constant cross-section, where the cross-section and initial data are defined as follows:
$a(x)$ =$=\begin{cases}\frac{1}{4},&0\leq x-\lfloor x\rfloor<\frac{1}{2},\\ \frac{3}{4},&\frac{1}{2}\leq x-\lfloor x\rfloor<1,\end{cases}$ (4.9a)
$\rho(x,0)$ $=0.3+\frac{1}{20}e^{-(x/8)^{2}},$ (4.9b)
$q(x,0)$ $=0.$ (4.9c)
The initial pulse divides into left- and right-going parts, each of which evolve into a train of traveling waves. In Figure 2, we show the right-going part, comparing the pseudospectral solution of the homogenized approximation with the finite volume solution of the variable cross-section isentropic gas system (2.2). We observe that the homogenized solution provides a good approximation at early times but, as expected, becomes less accurate as time progresses.
Figure 2: Comparison of the homogenized system (dashed line) and the variable-coefficient system (solid line) solutions, with the cross-section and initial data given by (4.9).
@@ 4.3.2 Smoothly-varying Cross-section
Next, we consider a scenario with a smoothly-varying cross-section, defined by the following initial data:
@@
$\rho(x,0)$ $=0.3+\frac{1}{12}e^{-(x/5)^{2}},$ (4.10b)
$q(x,0)$ $=0.$ (4.10c)
Figure 3: Comparison of the homogenized system (dashed line) and the variable coefficients system (solid line) solutions, with the cross-section and initial data given by (4.10).
The numerical solutions for this smooth cross-section are presented in Figure 3. Similar to the piecewise constant case discussed earlier, the accuracy of the homogenized approximation decreases over time.
@@ Appendix A Coefficients of the Homogenized Equations
$\vspace{5em}\hskip 17.50002ptC_{1}~{}$ $=\langle{a^{-1}\llbracket{a}\rrbracket}\rangle=-\langle{a\llbracket{a^{-1}}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{2}~{}$ $=\langle{a^{-1}\llbracket{\llbracket{a}\rrbracket}\rrbracket}\rangle=\langle{a\llbracket{\llbracket{a^{-1}}\rrbracket}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{3}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a\llbracket{a_{y}a^{-3}}\rrbracket}\rangle=\frac{-1}{(\underline{\rho}^{(0)})^{2}}\langle{a_{y}a^{-3}\llbracket{a}\rrbracket}\rangle=\frac{-1}{2(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-1}}\rangle-\langle{a}\rangle\langle{a^{-2}}\rangle\right),$
$\vspace{5em}\hskip 17.50002ptC_{4}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{\llbracket{a}\rrbracket}\rrbracket}\rangle=\frac{1}{(\underline{\rho}^{(0)})}\langle{a\llbracket{\llbracket{a^{-3}a_{y}}\rrbracket}\rrbracket}\rangle=\frac{-1}{2\underline{\rho}^{(0)}}\langle{a\llbracket{a^{-2}}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{5}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-1}\llbracket{a^{-3}a_{y}}\rrbracket}\rangle=\frac{-1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-3}a_{y}\llbracket{a^{-1}}\rrbracket}\rangle=\frac{-1}{2(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-3}}\rangle-\langle{a^{-1}}\rangle\langle{a^{-2}}\rangle\right),$
$\vspace{5em}\hskip 17.50002ptC_{6}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-3}a_{y}\llbracket{a\llbracket{a^{-3}a_{y}}\rrbracket}\rrbracket}\rangle=\frac{-1}{4(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-3}}\rangle+\langle{a}\rangle\langle{a^{-2}}\rangle^{2}-2\langle{a^{-1}}\rangle\langle{a^{-2}}\rangle\right),$
$\vspace{5em}\hskip 17.50002ptC_{7}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{a\llbracket{a^{-1}}\rrbracket}\rrbracket}\rangle=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-1}\llbracket{a\llbracket{a^{-3}a_{y}}\rrbracket}\rrbracket}\rangle=\frac{1}{2\underline{\rho}^{(0)}}\langle{a^{-2}}\rangle C_{1},$
$\vspace{5em}\hskip 17.50002ptC_{8}~{}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a\llbracket{a^{-3}a_{y}\llbracket{a}\rrbracket}\rrbracket}\rangle=\frac{-1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{a}\rrbracket^{2}}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{9}~{}$ $=\langle{a\llbracket{a^{-1}\llbracket{a}\rrbracket}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{10}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-1}}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{11}$ $=\langle{a^{-1}\llbracket{a\llbracket{a^{-1}}\rrbracket}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{12}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-3}}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{13}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a\llbracket{a^{-3}a_{y}}\rrbracket}\rangle=\frac{-1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{a}\rrbracket}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{14}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-1}}\rangle,$
$\vspace{5em}\hskip 17.50002ptC_{15}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a\llbracket{a^{-1}}\rrbracket}\rangle=\frac{-1}{(\underline{\rho}^{(0)})}\langle{a^{-1}\llbracket{a}\rrbracket}\rangle.$
@@
$\alpha_{1}$ $=\frac{-1}{\langle{a}\rangle},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha_{2}=\delta\frac{-C_{1}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha_{3}=\delta\frac{2C_{13}}{\langle{a^{-1}}\rangle\langle{a}\rangle},$
$\alpha_{4}$ $=\delta^{2}\left(\frac{4C_{13}^{2}+4C_{13}C_{14}}{\langle{a^{-1}}\rangle P^{\prime}\langle{a}\rangle^{2}}-\frac{C_{3}}{P^{\prime}\langle{a}\rangle^{2}}-\frac{C_{13}P^{\prime\prime}}{(P^{\prime})^{2}\langle{a}\rangle^{2}}\right),~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha_{5}=\delta^{2}\left(\frac{C_{9}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{3}}-\frac{C_{2}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}\right),$
$\alpha_{6}$ $=\delta^{2}\frac{-2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle},~{}~{}~{}~{}~{}~{}~{}~{}~{}\alpha_{7}=\delta^{2}\left(\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),~{}~{}~{}~{}~{}\alpha_{8}=\delta^{2}\left(-\frac{2C_{13}C_{1}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),$
@@
$\beta_{1}$ $=\frac{-P^{\prime}}{\langle{a^{-1}}\rangle},~{}~{}~{}~{}~{}~{}~{}\beta_{2}=\delta\left(\frac{-2C_{13}-2C_{14}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),~{}~{}~{}~{}~{}~{}\beta_{3}=\delta\frac{-P^{\prime\prime}}{\langle{a^{-1}}\rangle},$
$\beta_{4}$ $=\delta\frac{C_{1}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle},~{}~{}~{}~{}~{}~{}\beta_{5}=\delta^{2}\left(\frac{2C_{10}+2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),~{}~{}~{}~{}~{}~{}\beta_{6}=\delta^{2}\left(\frac{2C_{1}C_{13}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}-\frac{C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}\right),$
$\beta_{7}$ $=\delta^{2}\left(-\frac{2C_{1}C_{13}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{4C_{7}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),$
$\beta_{8}$ $=\delta^{2}\left(\frac{-3C_{5}+4C_{6}+C_{12}}{\langle{a^{-1}}\rangle^{2}}+\frac{4C_{13}^{2}+4C_{13}C_{14}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\right),$
$\beta_{9}$ $=\delta^{2}\left(\frac{2C_{7}P^{\prime}}{\langle{a^{-1}}\rangle^{3}}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{1}C_{13}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}\right),~{}~{}~{}~{}\beta_{10}=\delta^{2}\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle},~{}~{}~{}~{}\beta_{11}=\delta^{2}\left(\frac{C_{11}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}-\frac{C_{2}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\right).$

### v2

@@ 4.1 Numerical Discretization of Homogenized System
We discretize the full nonlinear homogenized equations (2.22), rewriting them first the form
@@
Here, $\mathcal{F}_{1}(\overline{\rho},\overline{q})$ and $\mathcal{F}_{2}(\overline{\rho},\overline{q})$ are the right-hand sides of the first and second equations in (2.22), respectively, excluding the xxtxxt-derivative terms. We solve this system using a Fourier pseudospectral collocation method in space and a third-order strong stability preserving Runge-Kutta method [24] in time.
@@ 4.3.1 Piecewise-Constant Cross-section
In this section, we consider flow in a pipe with piecewise-constant cross-section, where the cross-section and initial data are defined as follows. We conduct two different experiments corresponding to two choices of initial conditions.
$a(x)$ $=\begin{cases}\frac{1}{4},&0\leq x-\lfloor x\rfloor<\frac{1}{2},\\ \frac{3}{4},&\frac{1}{2}\leq x-\lfloor x\rfloor<1,\end{cases}$ (4.8a)
$\rho(x,0)$ $=1+\delta\,e^{-(x/8)^{2}},$ (4.8b)
$q(x,0)$ $=0,$ (4.8c)
Here, $0<\delta\ll 1$ is the Gaussian amplitude; we take $\delta=0.1$. In the first experiment, the Gaussian perturbation is superimposed on a constant background of $\rho=1$, so that the initial pulse divides into left- and right-going parts, each of which evolve into a train of traveling waves. In the second experiment, we instead take
$\rho(x,0)=0.3+\tfrac{1}{20}e^{-(x/8)^{2}},\qquad q(x,0)=0,$ (4.9)
which produces a similar splitting of the initial pulse, but with a smaller background state. Figures 3 and 4 show the right-going part for the first and second experiments, respectively. In both cases, we compare the pseudospectral solution of the homogenized approximation with the finite volume solution of the variable cross-section isentropic gas system (2.2). We observe that the homogenized solution provides a good approximation at early times but, as expected, becomes less accurate as time progresses.
Figure 3: First experiment ($\delta=0.1$): comparison of the homogenized system (dashed line) and the variable-coefficient system (solid line) solutions, with the cross-section and initial data given by (4.8).
Figure 4: Second experiment: comparison of the homogenized system (dashed line) and the variable-coefficient system (solid line) solutions, with the cross-section and initial data given by (4.8).
@@ 4.3.2 Smoothly-varying Cross-section
Next, we consider a scenario with a smoothly-varying cross-section. We conduct two different experiments corresponding to two choices of initial conditions.
@@
$\rho(x,0)$ $=1+\delta\,e^{-(x/5)^{2}},$ (4.10b)
$q(x,0)$ $=0,$ (4.10c)
Here, we use the same value of $\delta$ as in the piecewise-constant cross-section. In the first experiment, the Gaussian perturbation is superimposed on a constant background of $\rho=1$. In the second experiment, we instead take
$\rho(x,0)=0.3+\tfrac{1}{12}e^{-(x/5)^{2}},\qquad q(x,0)=0,$ (4.11)
which adds the Gaussian perturbation to a smaller background state. Figures 5 and 6 show the results for the first and second experiments, respectively. In both cases, we compare the pseudospectral solution of the homogenized approximation with the finite volume solution of the variable cross-section isentropic gas system (2.2). Similar to the piecewise constant case discussed earlier, the accuracy of the homogenized approximation decreases over time.
Figure 5: First experiment ($\delta=0.1$): comparison of the homogenized system (dashed line) and the variable coefficients system (solid line) solutions, with the cross-section and initial data given by (4.10).
Figure 6: Second experiment: comparison of the homogenized system (dashed line) and the variable coefficients system (solid line) solutions, with the cross-section and initial data given by (4.10).
@@ Appendix A Coefficients of the Homogenized Equations
$\vskip 50.00008pt\hskip 17.50002ptC_{1}$~ $=\langle{a^{-1}\llbracket{a}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{2}$~ $=\langle{a^{-1}\llbracket{\llbracket{a}\rrbracket}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{3}$~ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a\llbracket{a_{y}a^{-3}}\rrbracket}\rangle=\frac{-1}{2(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-1}}\rangle-\langle{a}\rangle\langle{a^{-2}}\rangle\right),$
$\vskip 50.00008pt\hskip 17.50002ptC_{4}$~ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{\llbracket{a}\rrbracket}\rrbracket}\rangle=\frac{-1}{2\underline{\rho}^{(0)}}\langle{a\llbracket{a^{-2}}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{5}$~ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-1}\llbracket{a^{-3}a_{y}}\rrbracket}\rangle=\frac{-1}{2(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-3}}\rangle-\langle{a^{-1}}\rangle\langle{a^{-2}}\rangle\right),$
$\vskip 50.00008pt\hskip 17.50002ptC_{6}$~ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-3}a_{y}\llbracket{a\llbracket{a^{-3}a_{y}}\rrbracket}\rrbracket}\rangle=\frac{-1}{4(\underline{\rho}^{(0)})^{2}}\left(\langle{a^{-3}}\rangle+\langle{a}\rangle\langle{a^{-2}}\rangle^{2}-2\langle{a^{-1}}\rangle\langle{a^{-2}}\rangle\right),$
$\vskip 50.00008pt\hskip 17.50002ptC_{7}$~ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-3}a_{y}\llbracket{a\llbracket{a^{-1}}\rrbracket}\rrbracket}\rangle=\frac{1}{2\underline{\rho}^{(0)}}\langle{a^{-2}}\rangle C_{1},$
$\vskip 50.00008pt\hskip 17.50002ptC_{8}$~ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a\llbracket{a^{-3}a_{y}\llbracket{a}\rrbracket}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{9}$~ $=\langle{a\llbracket{a^{-1}\llbracket{a}\rrbracket}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{10}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-1}}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{11}$ $=\langle{a^{-1}\llbracket{a\llbracket{a^{-1}}\rrbracket}\rrbracket}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{12}$ $=\frac{1}{(\underline{\rho}^{(0)})^{2}}\langle{a^{-3}}\rangle,$
$\vskip 50.00008pt\hskip 17.50002ptC_{13}$ $=\frac{1}{(\underline{\rho}^{(0)})}C_{3},$
$\vskip 50.00008pt\hskip 17.50002ptC_{14}$ $=\frac{1}{(\underline{\rho}^{(0)})}\langle{a^{-1}}\rangle.$
We make some simplifications using the identities
$\langle{f\llbracket{g}\rrbracket}\rangle$ $=-\langle{g\llbracket{f}\rrbracket}\rangle$
$\langle{f\llbracket{\llbracket{g}\rrbracket}\rrbracket}\rangle$ $=\langle{g\llbracket{\llbracket{f}\rrbracket}\rrbracket}\rangle$
$\llbracket{f^{-n}f^{\prime}}\rrbracket$ $=\frac{\langle{f^{1-n}}\rangle-f^{1-n}}{n-1}.$
@@
$\alpha_{1}$ $=\frac{-1}{\langle{a}\rangle},~~~~~~~~~~~~~~~~~\alpha_{2}=\frac{-C_{1}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}},~~~~~~~~~~~~~~~~~~~~~~~~~\alpha_{3}=\frac{2C_{13}}{\langle{a^{-1}}\rangle\langle{a}\rangle},$
$\alpha_{4}$ $=\frac{4C_{13}^{2}+4C_{13}C_{14}}{\langle{a^{-1}}\rangle P^{\prime}\langle{a}\rangle^{2}}-\frac{C_{3}}{P^{\prime}\langle{a}\rangle^{2}}-\frac{C_{13}P^{\prime\prime}}{(P^{\prime})^{2}\langle{a}\rangle^{2}},~~~~~~~~~\alpha_{5}=\frac{C_{9}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{3}}-\frac{C_{2}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}},$
$\alpha_{6}$ $=\frac{-2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle},~~~~~~~~~\alpha_{7}=\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle},~~~~~\alpha_{8}=-\frac{2C_{13}C_{1}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle},$
@@
$\beta_{1}$ $=\frac{-P^{\prime}}{\langle{a^{-1}}\rangle},~~~~~~~\beta_{2}=\frac{-2C_{13}-2C_{14}}{\langle{a^{-1}}\rangle\langle{a}\rangle},~~~~~~\beta_{3}=\frac{-P^{\prime\prime}}{\langle{a^{-1}}\rangle},$
$\beta_{4}$ $=\frac{C_{1}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle},~~~~~~\beta_{5}=\frac{2C_{10}+2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle},~~~~~~\beta_{6}=\frac{2C_{1}C_{13}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}-\frac{C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}},$
$\beta_{7}$ $=-\frac{2C_{1}C_{13}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{4C_{7}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle},$
$\beta_{8}$ $=\frac{-3C_{5}+4C_{6}+C_{12}}{\langle{a^{-1}}\rangle^{2}}+\frac{4C_{13}^{2}+4C_{13}C_{14}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle},$
$\beta_{9}$ $=\frac{2C_{7}P^{\prime}}{\langle{a^{-1}}\rangle^{3}}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{1}C_{13}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle},~~~~\beta_{10}=\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle},~~~~\beta_{11}=\frac{C_{11}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}-\frac{C_{2}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}.$